# Diffuse-CLoC: Guided Diffusion for Physics-based Character Look-ahead Control

### v2

@@ III-A Co-diffusing States and Actions
Architecture: Diffuse-CLoC uses a decoder-only transformer architecture, as shown in Fig. 1. In contrast to Diffuser [25], where a pair of state and action is combined into a single transformer token, our model takes the state and action input embeddings as separate tokens. Both the state and action tokens are concatenated from an MLP embedding and a sinusoidal embedding of their corresponding noise levels $\bm{k_{s}}$ and $\bm{k_{a}}$. We use a two-layer MLP for the state encoder and a linear layer for the action encoder. We then add learnable positional encodings to each token. The state and action embeddings over time are fed into a GPT-style decoder [39], followed by layer norm and the corresponding linear layers as output heads.
@@
Shorter Action Horizon: Predicting long state horizons, such as those extending to about a second, is essential for multi-modality and long-term planning required by certain tasks, but long-term action prediction is challenging due to increased variance [7]. To address this, we limit the action horizon to at most 16 steps and mask long-term future actions in the loss function, focusing on the initial actions only.
Emphasis Projection: To enhance state representation, we integrate emphasis projection [4] to emphasize the global states in the state space. Specifically, we define a projection matrix $\bm{P}=\bm{A}\bm{B}$, where $\bm{A}_{ij}\sim\mathcal{N}(0,1)$, and $\bm{B}$ a diagonal matrix with entries corresponding to the global states set to $c>1$ and others to one. Furthermore, unlike [4], sensitivity to local states is also preserved by concatenating projected and original states, i.e. $\bm{P}=[\bm{A}\bm{B}\ \ \mathbf{I}]$.
III-B Rolling Inference
@@
However, an issue in constant replanning is that planned trajectories could be inconsistent, causing oscillations. To address this issue, we use a rolling inference scheme. While a similar technique was introduced to generate infinite motions in kinematics space [9], we find it also improves the policy consistency during autoregressive execution. A FIFO buffer assigns noise levels based on the proximity of trajectory step to the current timestep. At each timestep, new Gaussian noise is added to the buffer while the oldest is removed. This approach leverages previous diffusion results for warmup, ensuring consistency and reducing diffusion steps. Since the the classifier guidance signal diminishes as $\bm{k}\rightarrow 0$, we perform rolling at a higher $\bm{k}$ for stronger guidance, while maintaining consistency and achieving speedup.
IV Applications
@@
IV-A Static Obstacle Avoidance
@@
$G_{\bm{\tau}}^{\text{obs}}(\bm{\tau})=\sum_{j}\sum_{t^{\prime}=t}^{t+H}\exp\big{(}-c\cdot\text{SDF}^{j}(\bm{s}_{t^{\prime}})\big{)},$ (3)
where $c$ determines the safe distance and $\text{SDF}^{j}(\cdot)$ is the Signed Distance Function from the $j$-th obstacle. For certain obstacles, such as a block on the ground, interaction (e.g., jumping onto it) may be desired rather than strict avoidance. In this case, we clip the SDF value outside the object to zero, penalizing only when the planned trajectory penetrates the obstacle. To achieve point-goal navigation while avoiding obstacles, the obstacle cost $G_{\bm{\tau}}^{\text{obs}}(\bm{\tau})$ can be combined with waypoint guidance, $G_{\bm{\tau}}^{\text{wp}}(\bm{\tau})=\sum_{t^{\prime}=t}^{t+H}\|P_{\text{root}}(\bm{s}_{t^{\prime}})-g\|^{2}$, where $P_{\text{root}}$ is the mapping from state representation to root position, and $g$ is the waypoint or goal location. We showcase different results using such costs in Fig. 1LABEL:fig:teaser, 6c, 6d, and 6e.
IV-B Dynamic Obstacle Avoidance
@@
$G_{\bm{\tau},i}^{\text{sa}}(\bm{\tau})=\sum_{j\neq i}\sum_{t^{\prime}=t}^{t+H}\exp\big{(}-c\cdot\|P_{\text{root}}(\bm{s}_{t^{\prime}}^{i})-P_{\text{root}}(\bm{s}_{t^{\prime}}^{j})\|^{2}\big{)},$ (4)
where $G_{\bm{\tau},i}^{\text{sa}}$ is the cost for the $i$-th character and $s_{t^{\prime}}^{j}$ is the state of the $j$-th character at timestep $t^{\prime}$. Fig. LABEL:fig:teaser shows an example of characters navigating through a forest of pillars while avoiding each other. This formulation allows characters to generate smooth trajectories that avoid collisions and pass each other seamlessly.
IV-C Task Space Control
@@
$G_{\bm{\tau}}^{\text{ts}}(\bm{\tau})=\sum_{t^{\prime}\in T}\|P_{x}(\bm{s}_{t^{\prime}})-g_{t^{\prime}}\|^{2},$ (5)
@@
Gamepad Controller Steering: Shown in Fig. LABEL:fig:teaser, setting the root’s velocity, heading, and height as the task space, with $g$ from gamepad inputs, enables real-time character steering.
IV-D Motion In-Betweening
In this task, the policy generates dense trajectories between a set of desired keyframes. With sparse keyframes, inpainting often struggles to differentiate clean from noisy states. While previous methods rely on extra masking to label these states [5], our approach, shown in Fig. LABEL:fig:teaser, can explicitly set the inpainted states to zero noise, naturally differentiating them from other noisy states.
$\bm{s}_{t}=\hat{\bm{s}}_{t},\quad\bm{{k_{s}}}_{t}=0,\quad\forall t\in T,$ (6)
@@ IV-D Motion In-Betweening
TABLE I: Benchmark for Kinematic-tracking baselines v.s. Diffuse-CLoC across various tasks.
Method Replan Interval Walk + Perturb Forest Jump Motion In-Between
@@ V Experiments and Results
V-A Data
We use a subset of AMASS [40], comprising 54 motions of walking, running, crawling, and jumping. To track these motions and collect state-action pairs, we employ PHC+ [41] as our motion tracking controller. For each motion, we collect on average 40 rollouts ($\approx 3.5$ hours). Following [7], we inject action noise during rollout to perturb the state and collect corrective actions. The state and action definitions are as follows:
@@ V-A Data
We use the SMPL skeleton from [17], with $J=23$ spherical joints. The overall state vector is 165-dimensional, consisting of Global and Local states defined as the following:
  • Global States: root states including position ($\mathbb{R}^{3}$), linear velocity ($\mathbb{R}^{3}$), and rotation ($\mathbb{R}^{3}$) represented as rotation vectors, as in [42]. These values are expressed relative to the character frame at the current timestep.
@@
V-B Experiment Setup
We train Diffuse-CLoC using an observation history of $N=4\ (\approx 0.13s)$ and predict future states for $H=32\ (\approx 1s)$. We limit the action horizon to $16$ steps. The model uses $20$ denoising steps, a transformer decoder with $6$ layers, $8$ heads, $512$-dimensional embedding, and attention dropout $p=0.3$.
@@ V-B1 Task & Metrics
[…]
@@
[…]
@@
[…]
V-B2 Baseline
The baseline we compare against represents the category of methods that tracks a diffused kinematic trajectory. As a hierarchical approach, the frequency of replanning is a key design consideration. Thus, we compare against Kin+PHC with different replan frequencies. We use the PHC+ tracking policy and separately train a kinematics-only diffusion model akin to the motion generation model in Guided-MDM [4].
V-B3 Ablation
We ablate over critical design choices, including prediction horizon for both state and action, as well as the number of diffusion steps. We also compare different attention schemes, including noncausal action attention, which removes action attention mask, and diffuser attention, which restricts attention to each token and its adjacent steps, following [25].
Figure 2: Rollouts in jump task for Kin+PHC (Left) vs. Diffuse-CLoC (Right). Kin+PHC suffers from artifacts in kinematics prediction and fails in agile motions, while Diffuse-CLoC remains robust and completes the task.
Figure 3: Rollouts in forest task for $0.5$s prediction (Left) vs. $1$s prediction (Right). Starting at the triangle, the character aims to reach the goal marked by a star. The short-horizon policy circles within the forest and fails to reach the waypoint, while the longer-horizon policy successfully navigates through the forest with diverse trajectories.
V-C Comparing Kinematics + Tracking
We compare Kin+PHC with Diffuse-CLoC across four tasks, finding that Diffuse-CLoC consistently outperforms all baseline variants (Table I). Kin+PHC exhibits a trade-off between motion quality and task success: slower replan frequencies improve motion smoothness (e.g., FID improves from $0.185$ at $1$-step replan to $0.038$ at $32$-step replan in the walk+perturb task) but fails to adapt to abrupt state changes. This results in a higher fall rate, which increases from $30\%$ at $1$-step to $44\%$ at $32$-step intervals. In contrast, Diffuse-CLoC maintains low FID ($0.074$) and reduces the fall rate to $16\%$, demonstrating robust performance with plausible motion quality.
@@ V-C Comparing Kinematics + Tracking
Unlike Kin+PHC, Diffuse-CLoC uses the joint distribution of states and actions, allowing it to fit over the entire horizon to generate dynamically feasible actions, even with partially corrupted kinematic inputs. By avoiding strict state tracking, Diffuse-CLoC reduces the infeasible motions from classifier guidance,Fig. 2.
V-D Effect of Prediction Horizon
The planning horizon significantly impacts the performance of Diffuse-CLoC (Table II), similar to its effect on conventional planning algorithms, where complex tasks require a sufficiently long planning horizon. With a shorter horizon of 16 timesteps ($\approx 0.5$s), the policy lacks foresight and struggles with forest navigation and complex maneuvers (squatting and jumping), leading to poor success rates. In contrast, extending the horizon to 32 timesteps ($\approx 1$s) dramatically improves performance, with forest navigation success at 96% and jumping tasks at 71%. Fig. 3 highlights this comparison visually: shorter horizons result in circular or stuck behaviors,while longer horizons enable successful task completion with diverse solution paths. However, further extending the horizon to $2$s reduces performance, likely because the increased variance in long-term states exerts excessive influence on the short-term predictions.
In contrast, a moderate action length improves action quality in agile motions like jumping. While one-step predictions cause jitter, longer predictions face higher future variance. Nevertheless, codiffusing states and actions allows it to leverage the lower state variance to improve the robustness of long action predictions, an advancement over prior works [8, 7]. In general, 32 state predictions combined with 16 action predictions yield optimal performance across our evaluation tasks.
TABLE II: Ablation study on horizon lengths, diffusion steps, attention styles, and rolling schemes. * marks the final model choice.
@@ V-D Effect of Prediction Horizon
$\checkmark$ $\times$ 58.06 79 14.75 76
$\times$ $\checkmark$ 58.06 96 10.62 43
$\times$ $\times$ 58.06 95 10.72 40
@@
V-E Effect of Different Attention Schemes
@@
Non-Causal Action Attention. Allowing actions to directly attend to future states (non-causal attention) decreases success rates (Table II). This decrease is due to two factors: (1) future states often include artifacts, and (2) due to the rolling scheme, future states tend to be noisier, making direct attention to future states unreliable.
@@
Diffuser’s Local Receptive Field. Diffuser [25] employs a limited receptive field, restricting attention to neighboring states and actions. This design prevents effective propagation of future information back to current actions, leading to poor coordination between state predictions and actions. Consequently, Diffuser struggles with task performance, as it cannot align actions with anticipated future conditions.
V-F Effect of Different Rolling Schemes

### v3

@@ 3.1. Co-diffusing States and Actions
Architecture: Diffuse-CLoC uses a decoder-only transformer architecture, as shown in Fig. 3. In contrast to Diffuser (Janner et al., 2022), where a pair of state and action is combined into a single transformer token, our model takes the state and action input embeddings as separate tokens. Both the state and action tokens are concatenated from an MLP embedding and a sinusoidal embedding of their corresponding noise levels $\bm{k_{s}}$ and $\bm{k_{a}}$. We use a two-layer MLP for the state encoder and a linear layer for the action encoder. We then add learnable positional encodings to each token. The state and action embeddings over time are fed into a GPT-style decoder (Radford, 2018), followed by layer norm and the corresponding linear layers as output heads.
@@
Shorter Action Horizon: Predicting long state horizons, such as those extending to about a second, is essential for multi-modality and long-term planning required by certain tasks, but long-term action prediction is challenging due to increased variance (Truong et al., 2024). To address this, we limit the action horizon to at most 16 steps and mask long-term future actions in the loss function, focusing on the initial actions only.
Emphasis Projection: To enhance state representation, we integrate emphasis projection (Karunratanakul et al., 2023) to emphasize the global states in the state space. Specifically, we define a projection matrix $\bm{P}=\bm{A}\bm{B}$, where $\bm{A}_{ij}\sim\mathcal{N}(0,1)$, and $\bm{B}$ a diagonal matrix with entries corresponding to the global states set to $c>1$ and others to one. Furthermore, unlike (Karunratanakul et al., 2023), sensitivity to local states is also preserved by concatenating projected and original states, i.e. $\bm{P}=[\bm{A}\bm{B}\ \ \mathbf{I}]$.
3.2. Rolling Inference
@@
However, an issue in constant replanning is that planned trajectories could be inconsistent, causing oscillations. To address this issue, we use a rolling inference scheme. While a similar technique was introduced to generate infinite motions in kinematics space (Zhang et al., 2024), we find it also improves the policy consistency during autoregressive execution. A FIFO buffer assigns noise levels based on the proximity of trajectory step to the current timestep. At each timestep, new Gaussian noise is added to the buffer while the oldest is removed. This approach leverages previous diffusion results for warmup, ensuring consistency and reducing diffusion steps. Since the the classifier guidance signal diminishes as $\bm{k}\rightarrow 0$, we perform rolling at a higher $\bm{k}$ for stronger guidance, while maintaining consistency and achieving speedup.
4. Applications
@@
4.1. Static Obstacle Avoidance
@@
(3) $G_{\bm{\tau}}^{\text{obs}}(\bm{\tau})=\sum_{j}\sum_{t^{\prime}=t}^{t+H}\exp\big{(}-c\cdot\text{SDF}^{j}(\bm{s}_{t^{\prime}})\big{)},$
where $c$ determines the safe distance and $\text{SDF}^{j}(\cdot)$ is the Signed Distance Function from the $j$-th obstacle. For certain obstacles, such as a block on the ground, interaction (e.g., jumping onto it) may be desired rather than strict avoidance. In this case, we clip the SDF value outside the object to zero, penalizing only when the planned trajectory penetrates the obstacle. To achieve point-goal navigation while avoiding obstacles, the obstacle cost $G_{\bm{\tau}}^{\text{obs}}(\bm{\tau})$ can be combined with waypoint guidance, $G_{\bm{\tau}}^{\text{wp}}(\bm{\tau})=\sum_{t^{\prime}=t}^{t+H}\|P_{\text{root}}(\bm{s}_{t^{\prime}})-g\|^{2}$, where $P_{\text{root}}$ is the mapping from state representation to root position, and $g$ is the waypoint or goal location. We showcase different results using such costs in Fig. 1, 4c, 4d, and 4e.
4.2. Dynamic Obstacle Avoidance
@@
(4) $G_{\bm{\tau},i}^{\text{sa}}(\bm{\tau})=\sum_{j\neq i}\sum_{t^{\prime}=t}^{t+H}\exp\big{(}-c\cdot\|P_{\text{root}}(\bm{s}_{t^{\prime}}^{i})-P_{\text{root}}(\bm{s}_{t^{\prime}}^{j})\|^{2}\big{)},$
where $G_{\bm{\tau},i}^{\text{sa}}$ is the cost for the $i$-th character and $s_{t^{\prime}}^{j}$ is the state of the $j$-th character at timestep $t^{\prime}$. Fig. 1 shows an example of characters navigating through a forest of pillars while avoiding each other. This formulation allows characters to generate smooth trajectories that avoid collisions and pass each other seamlessly.
4.3. Task Space Control
@@
(5) $G_{\bm{\tau}}^{\text{ts}}(\bm{\tau})=\sum_{t^{\prime}\in T}\|P_{x}(\bm{s}_{t^{\prime}})-g_{t^{\prime}}\|^{2},$
@@
Gamepad Controller Steering: Shown in Fig. 1, setting the root’s velocity, heading, and height as the task space, with $g$ from gamepad inputs, enables real-time character steering.
4.4. Motion In-Betweening
In this task, the policy generates dense trajectories between a set of desired keyframes. With sparse keyframes, inpainting often struggles to differentiate clean from noisy states. While previous methods rely on extra masking to label these states (Cohan et al., 2024), our approach, shown in Fig. 1, can explicitly set the inpainted states to zero noise, naturally differentiating them from other noisy states.
(6) $\bm{s}_{t}=\hat{\bm{s}}_{t},\quad\bm{{k_{s}}}_{t}=0,\quad\forall t\in T,$
@@ 4.4. Motion In-Betweening
Table 1. Benchmark for Kinematic-tracking baselines v.s. Diffuse-CLoC across various tasks.
Method
Replan
Interval
Walk + Perturb Forest Jump Motion In-Between
@@ 5. Experiments and Results
5.1. Data
We use a subset of AMASS (Mahmood et al., 2019), comprising 54 motions of walking, running, crawling, and jumping. To track these motions and collect state-action pairs, we employ PHC+ (Luo et al., 2024) as our motion tracking controller. For each motion, we collect on average 40 rollouts ($\approx 3.5$ hours). Following (Truong et al., 2024), we inject action noise during rollout to perturb the state and collect corrective actions. The state and action definitions are as follows:
@@ 5.1. Data
We use the SMPL skeleton from (Luo et al., 2023), with $J=23$ spherical joints. The overall state vector is 165-dimensional, consisting of Global and Local states defined as the following:
  • Global States: root states including position ($\mathbb{R}^{3}$), linear velocity ($\mathbb{R}^{3}$), and rotation ($\mathbb{R}^{3}$) represented as rotation vectors, as in (Zhang et al., 2023). These values are expressed relative to the character frame at the current timestep.
@@
5.2. Experiment Setup
We train Diffuse-CLoC using an observation history of $N=4(\approx 0.13s)$ and predict future states for $H=32(\approx 1s)$. We limit the action horizon to $16$ steps. The model employs a transformer decoder with $6$ layers, $8$ heads, $512$-dimensional embedding, totaling $19.95M$ parameters and requiring approximately 1 GB of GPU memory for inference. Training is conducted with $20$ denoising steps and an attention dropout rate of $p=0.3$
@@ 5.2.1. Task & Metrics
  • Walk-Perturb: We apply an instantaneous perturbation force sampled between 0 to $3000\,\text{N}$ every second during a 30s walk. We assess fall rate and motion quality using FID.
@@
  • Jump: We randomly place a $0.3$-$0.5\,\text{m}$ tall box obstacle $1$-$2\,\text{m}$ in front of the character. We measure success rate for clearing the obstacle without falling.
@@
For all tasks, a fall is determined if head height drops below $0.2\,\text{m}$.
5.2.2. Baseline
The baseline we compare against represents the category of methods that tracks a diffused kinematic trajectory. For the kinematics component, we train a diffusion model similar to the motion generation model in Guided-MDM (Karunratanakul et al., 2023), and for tracking we use PHC+ (Luo et al., 2024) . Since replanning frequency is a key design consideration in hierarchical approaches, we evaluate this baseline (Kin+PHC) across various replanning frequencies.
5.2.3. Ablation
We ablate over critical design choices, including prediction horizon for both state and action, as well as the number of diffusion steps. We also compare different attention schemes, including full attention, which enables unrestricted attention between all states and actions, and diffuser attention, which restricts attention to each token and its adjacent steps, following (Janner et al., 2022).
Figure 4. Downstream Tasks a) Task is to touch the red ball with the character’s right hand via classifier guidance. The red ball is re-located when the character achieves the task. b) Root path following via classifier guidance. We can further refine styles by constraining parameters like base velocity and heading. c) Walk and jump over two consecutive obstacles. d) Run and jump over obstacles. e) Jumping on different platforms using penetration cost and waypoint to each platform. f) A sequence of tasks in a single run, including route path following, platform jumping, and reaching a waypoint behind cylindrical barriers.
Figure 5. Rollouts in jump task for Kin+PHC (Left) vs. Diffuse-CLoC (Right). Kin+PHC suffers from artifacts in kinematics prediction and fails in agile motions, while Diffuse-CLoC remains robust and completes the task.
Figure 6. Rollouts in the forest task for $0.5$s prediction (Top), 11s prediction (Middle), and 22s prediction (Bottom). Starting at the triangle, the character aims to reach the goal marked by a star. The $0.5$s policy circles within the forest and fails to reach the waypoint, while the 11s policy successfully navigates through the forest with diverse trajectories. In contrast, the 22s policy sometimes overcommits to suboptimal future paths and collides with obstacles near the goal.
5.3. Comparing Kinematics + Tracking
We compare Kin+PHC with Diffuse-CLoC across four tasks, finding that Diffuse-CLoC consistently outperforms all baseline variants (Table 1). Kin+PHC exhibits a trade-off between motion quality and task success: slower replan frequencies improve motion smoothness (e.g., FID improves from $0.185$ at $1$-step replan to $0.038$ at $32$-step replan in the walk+perturb task) but fails to adapt to abrupt state changes. This results in a higher fall rate, which increases from $30\%$ at $1$-step to $44\%$ at $32$-step intervals. In contrast, Diffuse-CLoC maintains low FID ($0.074$) and reduces the fall rate to $16\%$, demonstrating robust performance with plausible motion quality.
@@ 5.3. Comparing Kinematics + Tracking
Unlike Kin+PHC, Diffuse-CLoC uses the joint distribution of states and actions, allowing it to fit over the entire horizon to generate dynamically feasible actions, even with partially corrupted kinematic inputs. By avoiding strict state tracking, Diffuse-CLoC reduces the infeasible motions from classifier guidance, as shown in Fig. 5.
Table 2. Ablation study on horizon lengths, diffusion steps, attention styles, and rolling schemes. * marks the final model choice.
@@
$\checkmark$
[…]
2613
58.06 79 14.75 76
[…]
2613
$\checkmark$ 58.06 96 10.62 43
[…]
2613
[…]
2613
58.06 95 10.72 40
@@
5.4. Effect of Prediction Horizon
The planning horizon significantly impacts the performance of Diffuse-CLoC (Table 2), similar to its effect on conventional planning algorithms, where complex tasks require a sufficiently long planning horizon. With a shorter horizon of 16 timesteps ($\approx 0.5$s), the policy lacks foresight and struggles with forest navigation and complex maneuvers (squatting and jumping), leading to poor success rates. In contrast, extending the horizon to 32 timesteps ($\approx 1$s) dramatically improves performance, with forest navigation success at 96% and jumping tasks at 71%. Fig. 6 highlights this comparison visually: a $0.5$s horizon results in circular or stuck behaviors, as the short-sighted policy fails to make progress through the forest, while a $1$s horizon enables successful task completion with diverse and adaptive solution paths that navigate around obstacles effectively. However, extending the horizon to $2$s reduces performance, as increased long-term prediction variance causes the policy to overcommit to suboptimal trajectories. As uncertainty decreases near obstacles, the policy recognizes the suboptimality but lacks flexibility to adjust.
In contrast, a moderate action length improves action quality in agile motions like jumping. While one-step predictions cause jitter, longer predictions face higher future variance. Nevertheless, codiffusing states and actions allows it to leverage the lower state variance to improve the robustness of long action predictions, an advancement over prior works (Huang et al., 2024a; Truong et al., 2024). In general, 32 state predictions combined with 16 action predictions yield optimal performance across our evaluation tasks.
Figure 7. Attention Maps for Obstacle Jumping. Each subplot shows how a state $s_{t}$ attends to tokens across the prediction horizon (x axis, 3-3 to 0 are observations) and over simulation time steps (y axis). First, the model learns localized attention to nearby states over time, indicated by the vertical color band around each state. Diagonal patterns at $s_{4}$, $s_{13}$, and $s_{29}$ indicate the model attends to the obstacle when predicted states encounter it with guidance cost and propagates this information backward for planning. Notably, future states also attend to past obstacle events to support planning. In contrast, $s_{0}$, the current state, consistently attends more strongly to past states than to future predictions, reducing the impact of artifacts from noisy future states.
5.5. Effect of Different Attention Schemes
@@
Full Attention. Allowing actions to directly attend to future states (full attention) decreases success rates (Table 2). This decrease is due to two factors: (1) future states often include artifacts, and (2) due to the rolling scheme, future states tend to be noisier, making direct attention to future states unreliable.
@@
Diffuser’s Local Receptive Field. Diffuser (Janner et al., 2022) employs a limited receptive field, restricting attention to neighboring states and actions. This design prevents effective propagation of future information back to current actions, leading to poor coordination between state predictions and actions. Consequently, Diffuser struggles with task performance, as it cannot align actions with anticipated future conditions.
5.6. Effect of Different Rolling Schemes